\newcommand{\etalchar}[1]{$^{#1}$}
\begin{thebibliography}{NAVB{\etalchar{+}}20}

\bibitem[Abr91]{a91}
Karl Abrahamson.
\newblock Time-space tradeoffs for algebraic problems on general sequential machines.
\newblock {\em Journal of Computer and System Sciences}, 43(2):269--289, 1991.

\bibitem[ADW{\etalchar{+}}24]{adw+24}
Josh Alman, Ran Duan, Virginia~Vassilevska Williams, Yinzhan Xu, Zixuan Xu, and Renfei Zhou.
\newblock More asymmetry yields faster matrix multiplication.
\newblock {\em arXiv preprint arXiv:2404.16349}, 2024.

\bibitem[AS23]{as23}
Josh Alman and Zhao Song.
\newblock Fast attention requires bounded entries.
\newblock {\em Advances in Neural Information Processing Systems}, 36, 2023.

\bibitem[AS24]{as24_arxiv}
Josh Alman and Zhao Song.
\newblock The fine-grained complexity of gradient computation for training large language models.
\newblock {\em arXiv preprint arXiv:2402.04497}, 2024.

\bibitem[BCS13]{bcs13}
Peter B{\"u}rgisser, Michael Clausen, and Mohammad~A Shokrollahi.
\newblock {\em Algebraic complexity theory}, volume 315.
\newblock Springer Science \& Business Media, 2013.

\bibitem[BGOFG20]{bgfg20}
Davis Blalock, Jose~Javier Gonzalez~Ortiz, Jonathan Frankle, and John Guttag.
\newblock What is the state of neural network pruning?
\newblock {\em Proceedings of machine learning and systems}, 2:129--146, 2020.

\bibitem[Bl{\"a}13]{bla13}
Markus Bl{\"a}ser.
\newblock Fast matrix multiplication.
\newblock {\em Theory of Computing}, pages 1--60, 2013.

\bibitem[BLSS20]{blss20}
Jan van~den Brand, Yin~Tat Lee, Aaron Sidford, and Zhao Song.
\newblock Solving tall dense linear programs in nearly linear time.
\newblock In {\em Proceedings of the 52nd Annual ACM SIGACT Symposium on Theory of Computing}, pages 775--788, 2020.

\bibitem[BMBE20]{bmbe20}
Brian Bartoldson, Ari Morcos, Adrian Barbu, and Gordon Erlebacher.
\newblock The generalization-stability tradeoff in neural network pruning.
\newblock {\em Advances in Neural Information Processing Systems}, 33:20852--20864, 2020.

\bibitem[BMR{\etalchar{+}}20]{bmr+20}
Tom Brown, Benjamin Mann, Nick Ryder, Melanie Subbiah, Jared~D Kaplan, Prafulla Dhariwal, Arvind Neelakantan, Pranav Shyam, Girish Sastry, Amanda Askell, Sandhini Agarwal, Ariel Herbert-Voss, Gretchen Krueger, Tom Henighan, Rewon Child, Aditya Ramesh, Daniel Ziegler, Jeffrey Wu, Clemens Winter, Chris Hesse, Mark Chen, Eric Sigler, Mateusz Litwin, Scott Gray, Benjamin Chess, Jack Clark, Christopher Berner, Sam McCandlish, Alec Radford, Ilya Sutskever, and Dario Amodei.
\newblock Language models are few-shot learners.
\newblock In H.~Larochelle, M.~Ranzato, R.~Hadsell, M.F. Balcan, and H.~Lin, editors, {\em Advances in Neural Information Processing Systems}, volume~33, pages 1877--1901. Curran Associates, Inc., 2020.

\bibitem[BPSW21]{bpsw21}
Jan van~den Brand, Binghui Peng, Zhao Song, and Omri Weinstein.
\newblock {Training (Overparametrized) Neural Networks in Near-Linear Time}.
\newblock In James~R. Lee, editor, {\em 12th Innovations in Theoretical Computer Science Conference (ITCS 2021)}, volume 185 of {\em Leibniz International Proceedings in Informatics (LIPIcs)}, pages 63:1--63:15, Dagstuhl, Germany, 2021. Schloss Dagstuhl -- Leibniz-Zentrum f{\"u}r Informatik.

\bibitem[Bra20]{bra20}
Jan van~den Brand.
\newblock A deterministic linear program solver in current matrix multiplication time.
\newblock In {\em Proceedings of the Fourteenth Annual ACM-SIAM Symposium on Discrete Algorithms}, pages 259--278. SIAM, 2020.

\bibitem[BSZ24]{bsz24}
Jan van~den Brand, Zhao Song, and Tianyi Zhou.
\newblock Algorithm and hardness for dynamic attention maintenance in large language models.
\newblock In {\em Forty-first International Conference on Machine Learning}, 2024.

\bibitem[CJD{\etalchar{+}}21]{cjd+21}
Tianyi Chen, Bo~Ji, Tianyu Ding, Biyi Fang, Guanyi Wang, Zhihui Zhu, Luming Liang, Yixin Shi, Sheng Yi, and Xiao Tu.
\newblock Only train once: A one-shot neural network training and pruning framework.
\newblock {\em Advances in Neural Information Processing Systems}, 34:19637--19651, 2021.

\bibitem[CLG{\etalchar{+}}24]{clg+24}
Tianle Cai, Yuhong Li, Zhengyang Geng, Hongwu Peng, Jason~D Lee, Deming Chen, and Tri Dao.
\newblock Medusa: Simple llm inference acceleration framework with multiple decoding heads.
\newblock {\em arXiv preprint arXiv:2401.10774}, 2024.

\bibitem[CLL{\etalchar{+}}24]{cll+24}
Bo~Chen, Xiaoyu Li, Yingyu Liang, Zhenmei Shi, and Zhao Song.
\newblock Bypassing the exponential dependency: Looped transformers efficiently learn in-context by multi-step gradient descent.
\newblock {\em arXiv preprint arXiv:2410.11268}, 2024.

\bibitem[CLP{\etalchar{+}}20]{clp+20}
Beidi Chen, Zichang Liu, Binghui Peng, Zhaozhuo Xu, Jonathan~Lingjie Li, Tri Dao, Zhao Song, Anshumali Shrivastava, and Christopher Re.
\newblock Mongoose: A learnable lsh framework for efficient neural network training.
\newblock In {\em International Conference on Learning Representations}, 2020.

\bibitem[CLS19]{cls19}
Michael~B Cohen, Yin~Tat Lee, and Zhao Song.
\newblock Solving linear programs in the current matrix multiplication time.
\newblock In {\em Proceedings of the 51st Annual ACM SIGACT Symposium on Theory of Computing}, pages 938--942, 2019.

\bibitem[CLS{\etalchar{+}}24]{cls+24}
Bo~Chen, Yingyu Liang, Zhizhou Sha, Zhenmei Shi, and Zhao Song.
\newblock Hsr-enhanced sparse attention acceleration.
\newblock {\em arXiv preprint arXiv:2410.10165}, 2024.

\bibitem[CTWC24]{ctwc24}
Ruisi Cai, Yuandong Tian, Zhangyang Wang, and Beidi Chen.
\newblock Lococo: Dropping in convolutions for long context compression.
\newblock {\em arXiv preprint arXiv:2406.05317}, 2024.

\bibitem[Dao23]{d23}
Tri Dao.
\newblock Flashattention-2: Faster attention with better parallelism and work partitioning.
\newblock {\em arXiv preprint arXiv:2307.08691}, 2023.

\bibitem[DFE{\etalchar{+}}22]{dfe+22}
Tri Dao, Dan Fu, Stefano Ermon, Atri Rudra, and Christopher R{\'e}.
\newblock Flashattention: Fast and memory-efficient exact attention with io-awareness.
\newblock {\em Advances in Neural Information Processing Systems}, 35:16344--16359, 2022.

\bibitem[DG24]{dg24}
Tri Dao and Albert Gu.
\newblock Transformers are ssms: Generalized models and efficient algorithms through structured state space duality.
\newblock {\em arXiv preprint arXiv:2405.21060}, 2024.

\bibitem[DMS24]{dms24}
Majid Daliri, Christopher Musco, and Ananda~Theertha Suresh.
\newblock Coupling without communication and drafter-invariant speculative decoding.
\newblock {\em arXiv preprint arXiv:2408.07978}, 2024.

\bibitem[DSW22]{dsw22}
Yichuan Deng, Zhao Song, and Omri Weinstein.
\newblock Discrepancy minimization in input-sparsity time.
\newblock {\em arXiv preprint arXiv:2210.12468}, 2022.

\bibitem[DYZ{\etalchar{+}}24]{dyz+24}
Harry Dong, Xinyu Yang, Zhenyu Zhang, Zhangyang Wang, Yuejie Chi, and Beidi Chen.
\newblock Get more with less: Synthesizing recurrence with kv cache compression for efficient llm inference.
\newblock {\em arXiv preprint arXiv:2402.09398}, 2024.

\bibitem[ESL{\etalchar{+}}24]{esl+24}
Mostafa Elhoushi, Akshat Shrivastava, Diana Liskovich, Basil Hosmer, Bram Wasti, Liangzhen Lai, Anas Mahmoud, Bilge Acun, Saurabh Agarwal, Ahmed Roman, et~al.
\newblock Layer skip: Enabling early exit inference and self-speculative decoding.
\newblock {\em arXiv preprint arXiv:2404.16710}, 2024.

\bibitem[FA22]{fa22}
Elias Frantar and Dan Alistarh.
\newblock Optimal brain compression: A framework for accurate post-training quantization and pruning.
\newblock {\em Advances in Neural Information Processing Systems}, 35:4475--4488, 2022.

\bibitem[FA23]{fa23}
Elias Frantar and Dan Alistarh.
\newblock Sparsegpt: Massive language models can be accurately pruned in one-shot.
\newblock In {\em International Conference on Machine Learning}, pages 10323--10337. PMLR, 2023.

\bibitem[FC18]{fc18}
Jonathan Frankle and Michael Carbin.
\newblock The lottery ticket hypothesis: Finding sparse, trainable neural networks.
\newblock {\em arXiv preprint arXiv:1803.03635}, 2018.

\bibitem[GD23]{gd23}
Albert Gu and Tri Dao.
\newblock Mamba: Linear-time sequence modeling with selective state spaces.
\newblock {\em arXiv preprint arXiv:2312.00752}, 2023.

\bibitem[GS22]{gs22}
Yuzhou Gu and Zhao Song.
\newblock A faster small treewidth sdp solver.
\newblock {\em arXiv preprint arXiv:2211.06033}, 2022.

\bibitem[GSWY23]{gswy23}
Yeqi Gao, Zhao Song, Weixin Wang, and Junze Yin.
\newblock A fast optimization view: Reformulating single layer attention in llm based on tensor and svm trick, and solving it in matrix multiplication time.
\newblock {\em arXiv preprint arXiv:2309.07418}, 2023.

\bibitem[GSZ23]{gsz23}
Yuzhou Gu, Zhao Song, and Lichen Zhang.
\newblock Faster algorithms for structured linear and kernel support vector machines, 2023.

\bibitem[GZL{\etalchar{+}}23]{gzl+23}
Suyu Ge, Yunan Zhang, Liyuan Liu, Minjia Zhang, Jiawei Han, and Jianfeng Gao.
\newblock Model tells you what to discard: Adaptive kv cache compression for llms.
\newblock {\em arXiv preprint arXiv:2310.01801}, 2023.

\bibitem[HABN{\etalchar{+}}21]{habn+21}
Torsten Hoefler, Dan Alistarh, Tal Ben-Nun, Nikoli Dryden, and Alexandra Peste.
\newblock Sparsity in deep learning: Pruning and growth for efficient inference and training in neural networks.
\newblock {\em Journal of Machine Learning Research}, 22(241):1--124, 2021.

\bibitem[HCI{\etalchar{+}}21]{hci+21}
Itay Hubara, Brian Chmiel, Moshe Island, Ron Banner, Joseph Naor, and Daniel Soudry.
\newblock Accelerated sparse neural training: A provable and efficient method to find n: m transposable masks.
\newblock {\em Advances in neural information processing systems}, 34:21099--21111, 2021.

\bibitem[HCL{\etalchar{+}}24]{hu2024outlier}
Jerry Yao-Chieh Hu, Pei-Hsuan Chang, Haozheng Luo, Hong-Yu Chen, Weijian Li, Wei-Po Wang, and Han Liu.
\newblock Outlier-efficient hopfield layers for large transformer-based models.
\newblock In {\em Forty-first International Conference on Machine Learning (ICML)}, 2024.

\bibitem[HCW{\etalchar{+}}24]{hu2024nonparametric}
Jerry Yao-Chieh Hu, Bo-Yu Chen, Dennis Wu, Feng Ruan, and Han Liu.
\newblock Nonparametric modern hopfield models.
\newblock {\em arXiv preprint arXiv:2404.03900}, 2024.

\bibitem[HJK{\etalchar{+}}24]{hjk+23}
Insu Han, Rajesh Jayaram, Amin Karbasi, Vahab Mirrokni, David Woodruff, and Amir Zandieh.
\newblock Hyperattention: Long-context attention in near-linear time.
\newblock In {\em The Twelfth International Conference on Learning Representations}, 2024.

\bibitem[HJS{\etalchar{+}}22]{hjs+22}
Baihe Huang, Shunhua Jiang, Zhao Song, Runzhou Tao, and Ruizhe Zhang.
\newblock Solving sdp faster: A robust ipm framework and efficient implementation.
\newblock In {\em 2022 IEEE 63rd Annual Symposium on Foundations of Computer Science (FOCS)}, pages 233--244. IEEE, 2022.

\bibitem[HLSL24]{hu2024computational}
Jerry Yao-Chieh Hu, Thomas Lin, Zhao Song, and Han Liu.
\newblock On computational limits of modern hopfield models: A fine-grained complexity analysis.
\newblock In {\em Forty-first International Conference on Machine Learning (ICML)}, 2024.

\bibitem[HNH{\etalchar{+}}21]{hnh+21}
Itay Hubara, Yury Nahshan, Yair Hanani, Ron Banner, and Daniel Soudry.
\newblock Accurate post training quantization with small calibration sets.
\newblock In {\em International Conference on Machine Learning}, pages 4466--4475. PMLR, 2021.

\bibitem[HPTD15]{hptd15}
Song Han, Jeff Pool, John Tran, and William Dally.
\newblock Learning both weights and connections for efficient neural network.
\newblock {\em Advances in neural information processing systems}, 28, 2015.

\bibitem[HS92]{hs92}
Babak Hassibi and David Stork.
\newblock Second order derivatives for network pruning: Optimal brain surgeon.
\newblock {\em Advances in neural information processing systems}, 5, 1992.

\bibitem[HWL24]{hwl24}
Jerry Yao-Chieh Hu, Dennis Wu, and Han Liu.
\newblock Provably optimal memory capacity for modern hopfield models: Tight analysis for transformer-compatible dense associative memories.
\newblock In {\em Advances in Neural Information Processing Systems (NeurIPS)}, volume~37, 2024.

\bibitem[HWSL24]{hwsl24}
Jerry Yao-Chieh Hu, Weimin Wu, Zhao Song, and Han Liu.
\newblock On statistical rates and provably efficient criteria of latent diffusion transformers (dits).
\newblock {\em arXiv preprint arXiv:2407.01079}, 2024.

\bibitem[HYW{\etalchar{+}}23]{hu2023sparse}
Jerry Yao-Chieh Hu, Donglin Yang, Dennis Wu, Chenwei Xu, Bo-Yu Chen, and Han Liu.
\newblock On sparse modern hopfield model.
\newblock In {\em Thirty-seventh Conference on Neural Information Processing Systems (NeurIPS)}, 2023.

\bibitem[JCR{\etalchar{+}}22]{jcr+22}
Tian Jin, Michael Carbin, Dan Roy, Jonathan Frankle, and Gintare~Karolina Dziugaite.
\newblock Pruning’s effect on generalization through the lens of training and regularization.
\newblock {\em Advances in Neural Information Processing Systems}, 35:37947--37961, 2022.

\bibitem[JKL{\etalchar{+}}20]{jkl+20}
Haotian Jiang, Tarun Kathuria, Yin~Tat Lee, Swati Padmanabhan, and Zhao Song.
\newblock A faster interior point method for semidefinite programming.
\newblock In {\em 2020 IEEE 61st annual symposium on foundations of computer science (FOCS)}, pages 910--918. IEEE, 2020.

\bibitem[JLSW20]{jlsw20}
Haotian Jiang, Yin~Tat Lee, Zhao Song, and Sam Chiu-wai Wong.
\newblock An improved cutting plane method for convex optimization, convex-concave games, and its applications.
\newblock In {\em Proceedings of the 52nd Annual ACM SIGACT Symposium on Theory of Computing}, pages 944--953, 2020.

\bibitem[JSWZ21]{jswz21}
Shunhua Jiang, Zhao Song, Omri Weinstein, and Hengjie Zhang.
\newblock A faster algorithm for solving general lps.
\newblock In {\em Proceedings of the 53rd Annual ACM SIGACT Symposium on Theory of Computing}, STOC 2021, page 823–832, New York, NY, USA, 2021. Association for Computing Machinery.

\bibitem[KKM{\etalchar{+}}22]{kks+22}
Woosuk Kwon, Sehoon Kim, Michael~W Mahoney, Joseph Hassoun, Kurt Keutzer, and Amir Gholami.
\newblock A fast post-training pruning framework for transformers.
\newblock {\em Advances in Neural Information Processing Systems}, 35:24101--24116, 2022.

\bibitem[KMZ23]{kmz23}
Praneeth Kacham, Vahab Mirrokni, and Peilin Zhong.
\newblock Polysketchformer: Fast transformers via sketches for polynomial kernels.
\newblock {\em arXiv preprint arXiv:2310.01655}, 2023.

\bibitem[LAT19]{lat19}
N~Lee, T~Ajanthan, and P~Torr.
\newblock Snip: single-shot network pruning based on connection sensitivity.
\newblock In {\em International Conference on Learning Representations}. Open Review, 2019.

\bibitem[LCH{\etalchar{+}}24]{lch+24}
Minghan Li, Xilun Chen, Ari Holtzman, Beidi Chen, Jimmy Lin, Wen-tau Yih, and Xi~Victoria Lin.
\newblock Nearest neighbor speculative decoding for llm generation and attribution.
\newblock {\em arXiv preprint arXiv:2405.19325}, 2024.

\bibitem[LDLG23]{ldlg23}
Yucheng Li, Bo~Dong, Chenghua Lin, and Frank Guerin.
\newblock Compressing context to enhance inference efficiency of large language models.
\newblock {\em arXiv preprint arXiv:2310.06201}, 2023.

\bibitem[LDS89]{lds89}
Yann LeCun, John Denker, and Sara Solla.
\newblock Optimal brain damage.
\newblock {\em Advances in neural information processing systems}, 2, 1989.

\bibitem[LG24]{lg24}
Fran{\c{c}}ois Le~Gall.
\newblock Faster rectangular matrix multiplication by combination loss analysis.
\newblock In {\em Proceedings of the 2024 Annual ACM-SIAM Symposium on Discrete Algorithms (SODA)}, pages 3765--3791. SIAM, 2024.

\bibitem[LGT{\etalchar{+}}21]{lgt+21}
Yuhang Li, Ruihao Gong, Xu~Tan, Yang Yang, Peng Hu, Qi~Zhang, Fengwei Yu, Wei Wang, and Shi Gu.
\newblock Brecq: Pushing the limit of post-training quantization by block reconstruction.
\newblock {\em arXiv preprint arXiv:2102.05426}, 2021.

\bibitem[LGU18]{lgu18}
Fran{\c{c}}ois Le~Gall and Florent Urrutia.
\newblock Improved rectangular matrix multiplication using powers of the coppersmith-winograd tensor.
\newblock In {\em Proceedings of the Twenty-Ninth Annual ACM-SIAM Symposium on Discrete Algorithms}, pages 1029--1046. SIAM, 2018.

\bibitem[LLS{\etalchar{+}}24a]{lls+24_io}
Xiaoyu Li, Yingyu Liang, Zhenmei Shi, Zhao Song, and Yufa Zhou.
\newblock Fine-grained attention i/o complexity: Comprehensive analysis for backward passes.
\newblock {\em arXiv preprint arXiv:2410.09397}, 2024.

\bibitem[LLS{\etalchar{+}}24b]{lls+24b}
Yingyu Liang, Heshan Liu, Zhenmei Shi, Zhao Song, and Junze Yin.
\newblock Conv-basis: A new paradigm for efficient attention inference and gradient computation in transformers.
\newblock {\em arXiv preprint arXiv:2405.05219}, 2024.

\bibitem[LLS{\etalchar{+}}24c]{lls+24_prune}
Yingyu Liang, Jiangxuan Long, Zhenmei Shi, Zhao Song, and Yufa Zhou.
\newblock Beyond linear approximations: A novel pruning approach for attention matrix.
\newblock {\em arXiv preprint arXiv:2410.11261}, 2024.

\bibitem[LM15]{lm15}
Shachar Lovett and Raghu Meka.
\newblock Constructive discrepancy minimization by walking on the edges.
\newblock {\em SIAM Journal on Computing}, 44(5):1573--1582, 2015.

\bibitem[LSS{\etalchar{+}}24]{lss+24}
Yingyu Liang, Zhizhou Sha, Zhenmei Shi, Zhao Song, and Yufa Zhou.
\newblock Multi-layer transformers gradient can be approximated in almost linear time.
\newblock {\em arXiv preprint arXiv:2408.13233}, 2024.

\bibitem[LSSY24]{lssy24}
Yingyu Liang, Zhenmei Shi, Zhao Song, and Chiwun Yang.
\newblock Toward infinite-long prefix in transformer.
\newblock {\em arXiv preprint arXiv:2406.14036}, 2024.

\bibitem[LSSZ24]{lssz24a}
Yingyu Liang, Zhenmei Shi, Zhao Song, and Yufa Zhou.
\newblock Tensor attention training: Provably efficient learning of higher-order transformers.
\newblock {\em arXiv preprint arXiv:2405.16411}, 2024.

\bibitem[LSZ19]{lsz19}
Yin~Tat Lee, Zhao Song, and Qiuyi Zhang.
\newblock Solving empirical risk minimization in the current matrix multiplication time.
\newblock In {\em Conference on Learning Theory}, pages 2140--2157. PMLR, 2019.

\bibitem[LSZ{\etalchar{+}}23]{lsz+23}
S.~Cliff Liu, Zhao Song, Hengjie Zhang, Lichen Zhang, and Tianyi Zhou.
\newblock {Space-Efficient Interior Point Method, with Applications to Linear Programming and Maximum Weight Bipartite Matching}.
\newblock In Kousha Etessami, Uriel Feige, and Gabriele Puppis, editors, {\em 50th International Colloquium on Automata, Languages, and Programming (ICALP 2023)}, volume 261 of {\em Leibniz International Proceedings in Informatics (LIPIcs)}, pages 88:1--88:14, Dagstuhl, Germany, 2023. Schloss Dagstuhl -- Leibniz-Zentrum f{\"u}r Informatik.

\bibitem[LTT{\etalchar{+}}24]{ltt+24}
Ji~Lin, Jiaming Tang, Haotian Tang, Shang Yang, Wei-Ming Chen, Wei-Chen Wang, Guangxuan Xiao, Xingyu Dang, Chuang Gan, and Song Han.
\newblock Awq: Activation-aware weight quantization for on-device llm compression and acceleration.
\newblock {\em Proceedings of Machine Learning and Systems}, 6:87--100, 2024.

\bibitem[LWD{\etalchar{+}}23]{lwd+23}
Zichang Liu, Jue Wang, Tri Dao, Tianyi Zhou, Binhang Yuan, Zhao Song, Anshumali Shrivastava, Ce~Zhang, Yuandong Tian, Christopher Re, et~al.
\newblock Deja vu: Contextual sparsity for efficient llms at inference time.
\newblock In {\em International Conference on Machine Learning}, pages 22137--22176. PMLR, 2023.

\bibitem[LZ20]{lz20}
Tianlin Liu and Friedemann Zenke.
\newblock Finding trainable sparse networks through neural tangent transfer.
\newblock In {\em International Conference on Machine Learning}, pages 6336--6347. PMLR, 2020.

\bibitem[MVK{\etalchar{+}}24]{mvk+24}
Jean Mercat, Igor Vasiljevic, Sedrick Keh, Kushal Arora, Achal Dave, Adrien Gaidon, and Thomas Kollar.
\newblock Linearizing large language models.
\newblock {\em arXiv preprint arXiv:2405.06640}, 2024.

\bibitem[NAVB{\etalchar{+}}20]{navb+20}
Markus Nagel, Rana~Ali Amjad, Mart Van~Baalen, Christos Louizos, and Tijmen Blankevoort.
\newblock Up or down? adaptive rounding for post-training quantization.
\newblock In {\em International Conference on Machine Learning}, pages 7197--7206. PMLR, 2020.

\bibitem[PMN{\etalchar{+}}23]{pmn+23}
Michael Poli, Stefano Massaroli, Eric Nguyen, Daniel~Y Fu, Tri Dao, Stephen Baccus, Yoshua Bengio, Stefano Ermon, and Christopher R{\'e}.
\newblock Hyena hierarchy: Towards larger convolutional language models.
\newblock In {\em International Conference on Machine Learning}, pages 28043--28078. PMLR, 2023.

\bibitem[QSZZ23]{qszz23}
Lianke Qin, Zhao Song, Lichen Zhang, and Danyang Zhuo.
\newblock An online and unified algorithm for projection matrix vector multiplication with application to empirical risk minimization.
\newblock In Francisco Ruiz, Jennifer Dy, and Jan-Willem van~de Meent, editors, {\em Proceedings of The 26th International Conference on Artificial Intelligence and Statistics}, volume 206 of {\em Proceedings of Machine Learning Research}, pages 101--156. PMLR, 25--27 Apr 2023.

\bibitem[SBZ{\etalchar{+}}24]{sbz+24}
Jay Shah, Ganesh Bikshandi, Ying Zhang, Vijay Thakkar, Pradeep Ramani, and Tri Dao.
\newblock Flashattention-3: Fast and accurate attention with asynchrony and low-precision.
\newblock {\em arXiv preprint arXiv:2407.08608}, 2024.

\bibitem[SCY{\etalchar{+}}24]{scy+24}
Hanshi Sun, Zhuoming Chen, Xinyu Yang, Yuandong Tian, and Beidi Chen.
\newblock Triforce: Lossless acceleration of long sequence generation with hierarchical speculative decoding.
\newblock {\em arXiv preprint arXiv:2404.11912}, 2024.

\bibitem[SLBK24]{slbk24}
Mingjie Sun, Zhuang Liu, Anna Bair, and J~Zico Kolter.
\newblock A simple and effective pruning approach for large language models.
\newblock In {\em The Twelfth International Conference on Learning Representations}, 2024.

\bibitem[SMN{\etalchar{+}}24]{smn+24}
Zhenmei Shi, Yifei Ming, Xuan-Phi Nguyen, Yingyu Liang, and Shafiq Joty.
\newblock Discovering the gems in early layers: Accelerating long-context llms with 1000x input token reduction.
\newblock {\em arXiv preprint arXiv:2409.17422}, 2024.

\bibitem[SSU18]{ssu18}
Mitchell Stern, Noam Shazeer, and Jakob Uszkoreit.
\newblock Blockwise parallel decoding for deep autoregressive models.
\newblock {\em Advances in Neural Information Processing Systems}, 31, 2018.

\bibitem[SY21]{sy21}
Zhao Song and Zheng Yu.
\newblock Oblivious sketching-based central path method for linear programming.
\newblock In {\em International Conference on Machine Learning}, pages 9835--9847. PMLR, 2021.

\bibitem[SY24]{sy24}
Barna Saha and Christopher Ye.
\newblock The i/o complexity of attention, or how optimal is flash attention?
\newblock {\em arXiv preprint arXiv:2402.07443}, 2024.

\bibitem[SYZ23]{syz23}
Zhao Song, Mingquan Ye, and Lichen Zhang.
\newblock Streaming semidefinite programs: $o(\sqrt{n})$ passes, small space and fast runtime, 2023.

\bibitem[SZZ24]{szz24}
Zhao Song, Lichen Zhang, and Ruizhe Zhang.
\newblock Training multi-layer over-parametrized neural network in subquadratic time.
\newblock In {\em 15th Innovations in Theoretical Computer Science Conference (ITCS 2024)}. Schloss-Dagstuhl-Leibniz Zentrum f{\"u}r Informatik, 2024.

\bibitem[TKYG20]{tkyg20}
Hidenori Tanaka, Daniel Kunin, Daniel~L Yamins, and Surya Ganguli.
\newblock Pruning neural networks without any data by iteratively conserving synaptic flow.
\newblock {\em Advances in neural information processing systems}, 33:6377--6389, 2020.

\bibitem[TZZ{\etalchar{+}}24]{tzz+24}
Jiaming Tang, Yilong Zhao, Kan Zhu, Guangxuan Xiao, Baris Kasikci, and Song Han.
\newblock Quest: Query-aware sparsity for efficient long-context llm inference.
\newblock {\em arXiv preprint arXiv:2406.10774}, 2024.

\bibitem[WHHL24]{wu2024uniform}
Dennis Wu, Jerry Yao-Chieh Hu, Teng-Yun Hsiao, and Han Liu.
\newblock Uniform memory retrieval with larger capacity for modern hopfield models.
\newblock In {\em Forty-first International Conference on Machine Learning (ICML)}, 2024.

\bibitem[WHL{\etalchar{+}}24]{wu2023stanhop}
Dennis Wu, Jerry Yao-Chieh Hu, Weijian Li, Bo-Yu Chen, and Han Liu.
\newblock {ST}anhop: Sparse tandem hopfield model for memory-enhanced time series prediction.
\newblock In {\em The Twelfth International Conference on Learning Representations (ICLR)}, 2024.

\bibitem[WXXZ24]{wxxz24}
Virginia~Vassilevska Williams, Yinzhan Xu, Zixuan Xu, and Renfei Zhou.
\newblock New bounds for matrix multiplication: from alpha to omega.
\newblock In {\em Proceedings of the 2024 Annual ACM-SIAM Symposium on Discrete Algorithms (SODA)}, pages 3792--3835. SIAM, 2024.

\bibitem[WZG19]{wzg19}
Chaoqi Wang, Guodong Zhang, and Roger Grosse.
\newblock Picking winning tickets before training by preserving gradient flow.
\newblock In {\em International Conference on Learning Representations}, 2019.

\bibitem[XHH{\etalchar{+}}24]{xu2024bishop}
Chenwei Xu, Yu-Chao Huang, Jerry Yao-Chieh Hu, Weijian Li, Ammar Gilani, Hsi-Sheng Goan, and Han Liu.
\newblock Bishop: Bi-directional cellular learning for tabular data with generalized sparse modern hopfield model.
\newblock In {\em Forty-first International Conference on Machine Learning (ICML)}, 2024.

\bibitem[XTC{\etalchar{+}}23]{xtc+23}
Guangxuan Xiao, Yuandong Tian, Beidi Chen, Song Han, and Mike Lewis.
\newblock Efficient streaming language models with attention sinks.
\newblock {\em arXiv preprint arXiv:2309.17453}, 2023.

\bibitem[Yes84]{y84}
Yaacov Yesha.
\newblock Time-space tradeoffs for matrix multiplication and the discrete fourier transform on any general sequential random-access computer.
\newblock {\em Journal of Computer and System Sciences}, 29(2):183--197, 1984.

\bibitem[YLG{\etalchar{+}}23]{ylg+23}
Hongru Yang, Yingbin Liang, Xiaojie Guo, Lingfei Wu, and Zhangyang Wang.
\newblock Pruning before training may improve generalization, provably.
\newblock {\em arXiv preprint arXiv:2301.00335}, 2023.

\bibitem[ZBKR24]{zbkr24}
Michael Zhang, Kush Bhatia, Hermann Kumbong, and Christopher R{\'e}.
\newblock The hedgehog \& the porcupine: Expressive linear attentions with softmax mimicry.
\newblock {\em arXiv preprint arXiv:2402.04347}, 2024.

\bibitem[ZDH24]{zdh24}
Amir Zandieh, Majid Daliri, and Insu Han.
\newblock Qjl: 1-bit quantized jl transform for kv cache quantization with zero overhead.
\newblock {\em arXiv preprint arXiv:2406.03482}, 2024.

\bibitem[ZHDK23]{zhdk23}
Amir Zandieh, Insu Han, Majid Daliri, and Amin Karbasi.
\newblock Kdeformer: Accelerating transformers via kernel density estimation.
\newblock In {\em International Conference on Machine Learning}, pages 40605--40623. PMLR, 2023.

\bibitem[ZHMK24]{zhmk24}
Amir Zandieh, Insu Han, Vahab Mirrokni, and Amin Karbasi.
\newblock Subgen: Token generation in sublinear time and memory.
\newblock {\em arXiv preprint arXiv:2402.06082}, 2024.

\bibitem[ZLC{\etalchar{+}}24]{zlc+24}
Zhenyu Zhang, Shiwei Liu, Runjin Chen, Bhavya Kailkhura, Beidi Chen, and Atlas Wang.
\newblock Q-hitter: A better token oracle for efficient llm inference via sparse-quantized kv cache.
\newblock {\em Proceedings of Machine Learning and Systems}, 6:381--394, 2024.

\bibitem[ZRG{\etalchar{+}}22]{zrg+22}
Susan Zhang, Stephen Roller, Naman Goyal, Mikel Artetxe, Moya Chen, Shuohui Chen, Christopher Dewan, Mona Diab, Xian Li, Xi~Victoria Lin, et~al.
\newblock Opt: Open pre-trained transformer language models.
\newblock {\em arXiv preprint arXiv:2205.01068}, 2022.

\bibitem[ZSZ{\etalchar{+}}24]{zsz+24}
Zhenyu Zhang, Ying Sheng, Tianyi Zhou, Tianlong Chen, Lianmin Zheng, Ruisi Cai, Zhao Song, Yuandong Tian, Christopher R{\'e}, Clark Barrett, et~al.
\newblock H2o: Heavy-hitter oracle for efficient generative inference of large language models.
\newblock {\em Advances in Neural Information Processing Systems}, 36, 2024.

\end{thebibliography}

\def\isarxiv{1} 

\ifdefined\isarxiv
\documentclass[11pt]{article}

\usepackage[numbers]{natbib}

\else

\documentclass{article}

\PassOptionsToPackage{numbers}{natbib}

\usepackage[final]{neurips_2024}

\usepackage[utf8]{inputenc} 
\usepackage[T1]{fontenc}    
\usepackage{hyperref}       
\usepackage{url}            
\usepackage{booktabs}       
\usepackage{amsfonts}       
\usepackage{nicefrac}       
\usepackage{microtype}      
\usepackage{xcolor}         

\fi

\usepackage{amsmath}
\usepackage{amsthm}
\usepackage{amssymb}
\usepackage{algorithm}
\usepackage{subfig}
\usepackage{algpseudocode}
\usepackage{graphicx}
\usepackage{grffile}
\usepackage{wrapfig,epsfig}
\usepackage{url}
\usepackage{xcolor}
\usepackage{epstopdf}

\allowdisplaybreaks

\ifdefined\isarxiv
\renewcommand*{\citet}{\cite}
\renewcommand*{\citep}{\cite}

\usepackage{tikz}
\usepackage{hyperref}  
\hypersetup{colorlinks=true,citecolor=blue,linkcolor=blue} 
\usetikzlibrary{arrows}
\usepackage[margin=1in]{geometry}

\else
\renewcommand*{\citet}{\cite}
\renewcommand*{\citep}{\cite}
\usepackage{microtype}
\usepackage{hyperref}
\definecolor{mydarkblue}{rgb}{0,0.08,0.45}
\hypersetup{colorlinks=true, citecolor=mydarkblue,linkcolor=mydarkblue}

\fi

\theoremstyle{plain}
\newtheorem{theorem}{Theorem}[section]
\newtheorem{lemma}[theorem]{Lemma}
\newtheorem{definition}[theorem]{Definition}

\newtheorem{fact}[theorem]{Fact}
\newtheorem{remark}[theorem]{Remark}

\newcommand{\wt}{\widetilde}

\newcommand{\R}{\mathbb{R}}

\newcommand{\Tmat}{{\cal T}_{\mathrm{mat}}}

\makeatletter
\newcommand*{\RN}[1]{\expandafter\@slowromancap\romannumeral #1@}
\makeatother

\usepackage{lineno}

\begin{document}

\ifdefined\isarxiv

\date{}

\title{A Tighter Complexity Analysis of SparseGPT}
\author{
Xiaoyu Li\thanks{\texttt{
xli216@stevens.edu}. Stevens Institute of Technology.}
\and
Yingyu Liang\thanks{\texttt{
yingyul@hku.hk}. The University of Hong Kong. \texttt{
yliang@cs.wisc.edu}. University of Wisconsin-Madison.} 
\and
Zhenmei Shi\thanks{\texttt{
zhmeishi@cs.wisc.edu}. University of Wisconsin-Madison.}
\and 
Zhao Song\thanks{\texttt{ magic.linuxkde@gmail.com}. The Simons Institute for the Theory of Computing at the University of California, Berkeley.}
}

\else

\title{A Tighter Complexity Analysis of SparseGPT} 

\author{%
  Xiaoyu Li \\
  Stevens Institute of Technology \\
  \texttt{xli216@stevens.edu} \\
    \And    
  Yingyu Liang \\
  The University of Hong Kong \\ 
  University of Wisconsin-Madison  \\
  \texttt{
yingyul@hku.hk, yliang@cs.wisc.edu} \\
  \And
  Zhenmei Shi \\
  University of Wisconsin-Madison \\
  \texttt{zhmeishi@cs.wisc.edu} \\
  \And
  Zhao Song \\
  The Simons Institute for the Theory of Computing \\ at the University of California, Berkeley \\
  \texttt{magic.linuxkde@gmail.com} \\
  \And
}

\maketitle 
\fi

\ifdefined\isarxiv
\begin{titlepage}
  \maketitle
  \begin{abstract}

In this work, we improved the analysis of the running time of SparseGPT [Frantar, Alistarh ICML 2023] from $O(d^{3})$ to $O(d^{\omega} + d^{2+a+o(1)} + d^{1+\omega(1,1,a)-a})$ for any $a \in [0, 1]$, where $\omega$ is the exponent of matrix multiplication. In particular, for the current $\omega \approx 2.371$ [Alman, Duan, Williams, Xu, Xu, Zhou 2024], our running time boils down to $O(d^{2.53})$. This running time is due to the analysis of the lazy update behavior in iterative maintenance problems such as [Deng, Song, Weinstein 2022; Brand, Song, Zhou ICML 2024].
  \end{abstract}
  \thispagestyle{empty}
\end{titlepage}

\newpage

\else

\begin{abstract}

\end{abstract}

\fi

\section{Introduction}\label{sec:intro} 

Large Language Models (LLMs) have been widely applied in many AI applications. 
Accelerating its inference speed is crucial to reduce the latency time during user usage. 
One of the recent brilliant works, SparseGPT~\citep{fa23}, uses calibration data to prune the parameters of GPT-family models~\citep{bmr+20, zrg+22}, by using the optimal brain damage technique~\citep{lds89, hs92}. 
Their algorithm (Algorithm~\ref{alg:sparse_gpt}) can prune at least 50\% parameters with structure patterns, while the perplexity increase is negligible. Thus, SparseGPT can reduce the running time and GPU memory usage while keeping high performance for LLMs' applications.  

SparseGPT~\citep{fa23} claims their main pruning algorithm (Algorithm~\ref{alg:sparse_gpt}) takes running time $O(d^3)$ where $d$ is the model's hidden feature dimensions. In our work, we improved the running time analysis to get a tighter running time complexity bound, i.e., $O(d^{2.53})$, for Algorithm~\ref{alg:sparse_gpt}. Formally, we state our main results as follows.

\begin{theorem}[Main result (Restatement of Theorem~\ref{thm:main})]\label{thm:main_informal}
    Let lazy update block size $B = d^a$ for any $a \in [0, 1]$. Then Procedure \textsc{SparseGPT} in Algorithm~\ref{alg:sparse_gpt} achieves the running time
    \begin{align*}
        O(d^{\omega} + d^{2+a+o(1)} + d^{1+\omega(1,1,a)-a}).
    \end{align*}
\end{theorem}
For the current $\omega \approx 2.371$, according to the Table~1 in~\citet{adw+24}, we should choose $a\approx 0.5275$ in order to balance the terms $d^{2+a+o(1)}$ and $d^{1+\omega(1,1,a)-a}$. Then the running time boils down to $O(d^{2.53})$, which is better than $O(d^3)$ as claimed in~\citet{fa23}.

The key technique in our analysis is to compute the complexity of SparseGPT by using lazy update. 
It comes from an interesting fact of fast rectangular matrix multiplication: the time complexity of multiplying a $d \times d$ matrix by a
$d \times 1$ matrix is the same as the times complexity of multiplying a $d \times d$ matrix by a $d \times d^{a}$ matrix for any nonnegative $a \leq \alpha$ where $\alpha$ is the dual exponent of matrix multiplication and currently $\alpha \approx 0.321$~\citep{wxxz24, lg24, lgu18}. 
The lazy update has been widely used to speed up the training of neural networks~\citep{clp+20, bpsw21, szz24} and maintain dynamic attention in large language models~\citep{bsz24}.
The high-level intuition of lazy update is that during iterative one-rank updates for a $d \times d$ matrix, 
we use an external matrix $d\times d^a$ to save the input for updates while 
we conduct fast matrix multiplication only when $d^a$ updates happen. Based on the lazy update idea, we achieve our tighter complexity bound.  
\section{Related Work}\label{sec:related}

\subsection{Model Acceleration}
Model acceleration is critical and urgently needed in many practical applications. There are many techniques for model acceleration. One line of work is to change model architecture to support fast inference, e.g., Mamba~\citep{gd23,dg24}, PolySketchFormer~\citep{kmz23}, Linearizing Transformers~\citep{zhdk23,zbkr24,mvk+24}, Hopfield Model~\citep{hu2024nonparametric,hu2024outlier,wu2024uniform,xu2024bishop,hu2024computational,wu2023stanhop,hu2023sparse,hwl24} and so on. 
Another line of work is accelerating model computation on system level, e.g., FlashAttetnion~\citep{dfe+22,d23,sbz+24,sy24,lls+24_io}, parallel decoding~\citep{ssu18}, quantization~\citep{zlc+24,tzz+24,ltt+24, zdh24} and so on. 
To accelerate LLMs training and inference, there is a line of work to approximate attention matrix computation~\citep{as23,as24_arxiv,hjk+23,zhmk24,lssz24a,pmn+23,ctwc24,lssy24,lls+24b, gswy23,lssy24,lss+24,cll+24,hwsl24,cls+24} in almost linear time. 
Some specific technique is developed to accelerate LLMs generation, e.g., KV-Cache compression~\citep{gzl+23,ldlg23,xtc+23,zsz+24,lwd+23,dyz+24,scy+24,smn+24} and speculative decodings~\citep{clg+24,lch+24, scy+24, esl+24, dms24}.  

\subsection{Model Pruning}
Pruning is a technique aimed at reducing the number of weights in a neural network by selectively removing certain neurons~\citep{lds89, hs92}. This approach has gained considerable attention in recent years as a method to enhance the efficiency and scalability of deep learning models~\citep{hptd15, fc18, lat19, wzg19, bgfg20, bmbe20, lz20, tkyg20, cjd+21, habn+21, hci+21, jcr+22, fa22, fa23, slbk24}.
Pruning can be categorized based on its stage in the training process: pre-training pruning and post-training pruning. Pre-training pruning involves pruning the network at initialization. \citet{fc18} demonstrated that a neural network pruned at initialization could be trained to achieve performance comparable to a dense model, a phenomenon referred to as the lottery ticket hypothesis. This discovery motivated a line of research focused on developing methods to reduce the computational cost of pruning neural networks at initialization~\citep{lat19, wzg19, lz20, tkyg20, cjd+21}. More recently, \citet{ylg+23} provided theoretical evidence that pre-training pruning can enhance a model's generalization ability.
Post-training pruning, initially popularized through its application in quantization~\citet{navb+20, hnh+21, lgt+21, slbk24}, was later extended to pruning by \citet{hci+21, fa22, kks+22, lls+24_prune}. Post-training pruning aims to compress a well-optimized model using a small set of calibration data. This process typically involves layer-wise pruning of the neural network. Notably, \citet{hci+21} proposed a provable and efficient method for compressing a model by ``stitching together'' individually compressed layers.

\subsection{Lazy Update}

In recent years, the lazy update idea used in iterative maintenance problems has emerged as a crucial technique to effectively and efficiently solve various optimization problems, including linear programming~\citep{cls19, bra20, blss20, jswz21, sy21, lsz+23}, semi-definite programming~\citep{jkl+20, hjs+22, gs22, syz23}, empirical risk minimization~\citep{lsz19, gsz23, qszz23}, cutting plane methods~\citep{jlsw20}, neural network training~\citep{clp+20, bpsw21, szz24}, discrepancy minimization~\citep{dsw22}, dynamic attention problems~\citep{bsz24} and so on. 

We highlight several previous papers that share similar running time complexity. For comparison, we assume the input size of these problems is dominated by a parameter $d$. In the line of developments of fast linear program solvers, \citet{cls19} first introduced the lazy update idea to develop efficient dynamic inverse structures for implementing interior point methods in solving linear programming problems, and it achieves the running time $O^*(d^\omega + d^{2.5-a/2} + d^{1.5+a})$. \citet{sy21} uses a different method to achieve linear programming in the same complexity $O^*(d^\omega + d^{2.5-a/2} + d^{1.5+a})$ by incorporating the sparse sketching matrix to speed up the online matrix-vector multiplication. \citet{jswz21} improves the linear programming to $O^*(d^\omega + d^{2.5 - a/2} + d^{1.5 + a - \wt{a}/2} + d^{0.5+a+(\omega-1)\wt{a}})$ time where $\wt{a} \in [0, \alpha a]$ by designing a two-level lazy update framework for efficiently maintaining a projection matrix.
Later, \citet{lsz19} generalizes the algorithm for linear programming to empirical risk minimization, which also takes $O^*(d^\omega + d^{2.5-a/2} + d^{1.5+a})$ time. Different from~\citet{cls19}, which employs a stochastic central path method that updates weights using a random sparse vector, \citet{lsz19} introduces a robust deterministic central path method. Additionally, \citet{lsz19} proposes an efficient data structure capable of handling updates efficiently even when the weight update vector is dense. Further, \citet{qszz23} uses a different method to achieve empirical risk minimization in the same complexity $O^*(d^\omega + d^{2.5-a/2} + d^{1.5+a})$ via online projection matrix-vector multiplication. 

There are numerous other applications of the lazy update technique. For instance, \citet{dsw22} employs this approach to solve a subroutine of the discrepancy minimization problem in $O^*(d^\omega + d^{2+a} + d^{1 + \omega(1,1,a) - a})$. This is achieved by designing a data structure that efficiently implements the iterative Edge-Walk partial-coloring algorithm proposed by \citet{lm15}, utilizing the lazy update concept. More recently, \citet{bsz24} demonstrates how the lazy update concept can be applied to a dynamic attention problem, achieving an amortized update time of $O(d^{\omega(1,1,a)-a})$.
\section{Preliminary}\label{sec:preli}
\subsection{Notations.}
For a matrix $A$, we denote its transpose by $A^\top$. We use $I_{d\times d}$ to denote a $d\times d$ identity matrix. We use $\mathbf{1}_{n \times d}$ to denote an $n \times d$ matrix where all entries are ones and $\mathbf{0}_{n \times d}$ to denote an $n \times d$ matrix where all entries are zeros. For a matrix $A$, the notation $A_{[i_1, i_2], [j_1, j_2]}$ refers to the submatrix of $A$ corresponding to the rows from $i_1$ to $i_2$ (including $i_1$ and $i_2$) and columns from $j_1$ to $j_2$ (including $j_1$ and $j_2$). When we write $A_{*, [j_1, j_2]}$, it denotes the submatrix of $A$ that includes all rows and restricts the columns to those from $j_1$ to $j_2$. For two matrices $A, B$ of the same size, we denote the Hadamard product of $A$ and $B$ by $A \circ B$. We use
$O^*(f(d))$ to hide $f(d)^{o(1)}$ factor.

\subsection{Definitions and Facts} 
We now introduce some key definitions and key facts.
\begin{definition}
    For three positive integers $d_1, d_2, d_3$, we use $\Tmat(d_1, d_2, d_3)$ to denote the time of multiplying a $d_1 \times d_2$ matrix with a $d_2 \times d_3$ matrix. 
\end{definition}

The following fact shows that the order of $d_1, d_2, d_3$ only results in a constant factor difference. 

\begin{fact}[\citet{bcs13,bla13}]
    It holds that
    \begin{align*}
        \Tmat(d_1, d_2, d_3) = O(\Tmat(d_1, d_3, d_2)) = O(\Tmat(d_2,d_1,d_3)).
    \end{align*}
\end{fact}
We provide a definition of $\omega(\cdot,\cdot,\cdot)$ here. 
\begin{definition}
\label{def:omega_alpha}
For $a,b,c>0$, we use $d^{w(a,b,c)}$ to denote the time complexity of multiplying a $d^a \times d^b$ matrix with a $d^b \times d^c$ matrix. We define $\omega := \omega(1,1,1)$ as the exponent of matrix multiplication. We use $\alpha$ to denote the dual exponent of matrix multiplication, which is the largest value such that $\omega(1, \alpha, 1) = 2 + o(1)$. 
\end{definition}

In other words, $\omega$ means that multiplying two $d \times d$ matrices require time $O(d^\omega)$, and $\alpha$ is the largest number such that we can multiply a $d \times d^\alpha$ matrix with a $d^\alpha \times d$ in the subquadratic time.

\begin{lemma}[\citet{adw+24, wxxz24, lg24}]
    Currently, we have $\omega \approx 2.371$ and $\alpha \approx 0.321$. 
\end{lemma}

\begin{algorithm}[!ht]\caption{The SparseGPT algorithm (Algorithm~1 in~\citet{fa23}).}
\label{alg:sparse_gpt}
\begin{algorithmic}[1]
\Procedure{SparseGPT}{$p\in[0,1]$, $W \in \R^{d \times d}$, $X \in \R^{d \times d}$, $B \in \mathbb N_+$, $B_s \in \mathbb N_+$, $\lambda > 0$}
\State \Comment{Pruning ration $p \in [0,1]$}
\State \Comment{Weight matrix $W \in \R^{d \times d}$}
\State \Comment{Input feature matrix $X \in \R^{d \times d}$}
\State \Comment{Lazy update block size $B \in \mathbb N_+$, $B=d^a$ for any $a\in[0,1]$}
\State \Comment{Adaptive mask size $B_s \in  \mathbb N_+$}
\State \Comment{Regularization parameter $\lambda > 0$}
\State $M, ~ E \gets {\bf 1}_{d \times d}, ~ {\bf 0}_{d \times B}$ \label{line:init_m} \Comment{$O(d^2)$}
\State $\wt{H} \gets (XX^\top + \lambda I_{d \times d})^{-1}$  \Comment{$O(d^\omega)$ by Lemma~\ref{lem:run_time_h}} \label{line:init_h} 
\For{$i = 0, B, 2B, \ldots, \lfloor \frac{d}{B} \rfloor B$}
\For{$j = i + 1, \ldots, i+B$}
\If{$j \bmod B_s = 0$} \label{line:check_j} \Comment{$O(d)$}
\State $M_{*,[j, j+B_s]} \gets \textsc{MaskSelect}(p, W_{*, [j, j+B_s]}, \wt{H}, j - 1)$ \label{line:update_m}
\Comment{$O(d^2 \log d)$ by Lemma~\ref{lem:run_time_mask}}
\EndIf
\State $E_{*,j-i} \gets (\mathbf{1}_{d \times 1}-M_{*,j})\circ W_{*,j}/\wt{H}_{j, j}$ \label{line:update_e_2} \Comment{$O(d^2)$}
\State $W_{*,[j, i+B]} \gets W_{*, [j, i+B]} - E_{*,j-i} \wt{H}_{j,[j, i+B]}$ \label{line:update_w_inner}
\Comment{$O(d^{2+a})$ by Lemma~\ref{lem:run_time_2}}
\EndFor
\State $W_{*,[i+B, d]} \gets W_{*, [i+B,d]} - E \wt{H}_{[i, i+B],[i+B, d]}$ \label{line:update_w_outter}
\Comment{$O(d^{1+\omega(1,1,a) - a})$ by Lemma~\ref{lem:run_time_1}}
\EndFor
\State $W \gets W \circ M$ \Comment{$O(d^2)$} \label{line:final_w}
\EndProcedure
\State
\Procedure{MaskSelect}{$p \in [0,1]$, $W' \in \R^{d \times r}$, $\wt{H} \in \R^{d \times d}, s \in \mathbb{N}_+$} 
\State \Comment{Sub-weight matrix $W' \in \R^{d \times r}$; Inverse of Hessian matrix $\wt{H} \in \R^{d \times d}$}
\State \Comment{Index $s \in \mathbb{N}_+$, recording the position of $W'$ in $W$}
\State $M' \gets \mathbf{0}_{d \times r}$ \label{line:sub_init_m} \Comment{$O(dr)$}
\For{$k = 1, \ldots, r$}
\State $w \gets W'_{*, k} $ \label{line:sub_init_w} \Comment{$w\in\R^d$, $O(dr)$}
\State $w \gets (w \circ w)/ (\wt{H}_{s+k, s+k})^2$ \label{line:sub_update_w} \Comment{$O(dr)$}
\State $J \gets $ \label{line:sub_update_i} indices of top $(1-p)d$ largest entries of $w$ \Comment{$O(r \cdot d\log d)$ by sorting}
\For{$j \in J$}
\State $M'_{k, j} \gets 1$ \label{line:sub_update_m_ki} \Comment{$O(dr)$}
\EndFor
\EndFor
\State \Return{$M'$}
\EndProcedure
\end{algorithmic}
\end{algorithm}

\section{Main Results}\label{sec:analysis}

In this section, we present our principal findings. Our analysis demonstrates that SparseGPT (refer to Algorithm~\ref{alg:sparse_gpt}) attains the desired running time of $ O(d^{\omega} + d^{2+a} + d^{1+\omega(1,1,a)-a}) $ (Theorem~\ref{thm:main}). For the sake of clarity, we assume that both the weight matrix $ W $ and the input feature matrix $ X $ are of dimensions $ \mathbb{R}^{d \times d} $. However, our analysis remains valid for more general cases where $ W \in \mathbb{R}^{n \times d} $ and $ X \in \mathbb{R}^{d \times N} $, provided that $ n = O(d) $ and $N = O(d) $.

\begin{theorem}[Main result]\label{thm:main}
    Let lazy update block size $B = d^a$ for any $a \in [0, 1]$. Then Procedure \textsc{SparseGPT} in Algorithm~\ref{alg:sparse_gpt} achieves the running time
    \begin{align*}
        O^*(d^{\omega} + d^{2+a} + d^{1+\omega(1,1,a)-a}).
    \end{align*}
\end{theorem}
\begin{proof}
    We split the analysis of running time in the following
    \begin{itemize}
        \item Line~\ref{line:init_m} takes time $O(d^2)$ to initiate $M$ and takes time $O(d^{1+a})$ to initiate $E$.
        \item By Lemma~\ref{lem:run_time_h}, Line~\ref{line:init_h} takes time $O(d^\omega)$.
        \item In each iteration, Line~\ref{line:check_j} takes time $O(1)$ to check if $j \bmod B_s = 0$. Since there are $d$ iterations, the total time is $O(d)$.
        \item By Lemma~\ref{lem:run_time_mask}, Line~\ref{line:update_m} takes time $O(d^2\log d) = O(d^{2+o(1)})$ over all iterations.
        \item In each iteration, Line~\ref{line:update_e_2} takes time $O(d)$ to compute $W_{*,j}/[\wt{H}]_{jj}^2$ and takes time $O(d)$ to compute $1_{d \times 1}-M_{*,j}$ and then takes time $O(d)$ to compute the Hadamard product of them. Since there are $d$ iterations, the total time is $O(d^2)$.
        \item By Lemma~\ref{lem:run_time_2}, Line~\ref{line:update_w_inner} takes time $O(d^{2+a})$ over all iterations.
        \item By Lemma~\ref{lem:run_time_1}, Line~\ref{line:update_w_outter} takes time $O(d^{1+\omega(1,1,a) - a})$ over all iterations.
        \item Line~\ref{line:final_w} takes time $O(d^2)$ to compute the Hadamard product of two $d\times d$ matrices.
    \end{itemize}
    Summing up,  Algorithm~\ref{alg:sparse_gpt} runs in time
    \begin{align*}
        O(\underbrace{d^2}_{\mathrm{Line}~\ref{line:init_m}} + \underbrace{d^\omega}_{\mathrm{Line}~\ref{line:init_h}} + \underbrace{d}_{\mathrm{Line}~\ref{line:check_j}} + \underbrace{d^{2+o(1)}}_{\mathrm{Line}~\ref{line:update_m}} + \underbrace{d^2}_{\mathrm{Line}~\ref{line:update_e_2}} + \underbrace{d^{2+a}}_{\mathrm{Line}~\ref{line:update_w_inner}} + \underbrace{d^{1+\omega(1,1,a)-a}}_{\mathrm{Line}~\ref{line:update_w_outter}} + \underbrace{d^2}_{\mathrm{Line}~\ref{line:final_w}}).
    \end{align*}
    Absorbing the same terms and lower order terms, we get our target running time
   \begin{align*}
        O(d^{\omega} + d^{2+a+o(1)} + d^{1+\omega(1,1,a)-a}).
    \end{align*}
    The proof is complete.
\end{proof}

\begin{remark}
    The $O(d^2)$ space complexity is sufficient to implement fast matrix multiplication, which is equivalent to that of standard matrix multiplication. In fact, all algorithms derived from Strassen's original approach exhibit a $\Theta(d^2)$ space complexity. For further details, we refer readers to \cite{y84, a91}.
\end{remark}

\section{Detailed Complexity Analysis}

From now on, we let the lazy update block size $ B = d^a $ for any $ a \in [0, 1]$ without defining it in the statement of lemmas. We first analyze the running time of the subroutine Procedure \textsc{MaskSelect} in Algorithm~\ref{alg:sparse_gpt}.

\begin{lemma}\label{lem:run_time_mask}
The running time of Procedure \textsc{MaskSelect} in Algorithm~\ref{alg:sparse_gpt} is $O(rd\log d)$. Hence the running time of Line~\ref{line:update_m} in Algorithm~\ref{alg:sparse_gpt} over all iterations is $O(d^2 \log d)$.
\end{lemma}
\begin{proof}
    We first analyze the running time of Procedure \textsc{MaskSelect}.
    We split the analysis of running time in the following
    \begin{itemize}
        \item Line~\ref{line:sub_init_m} takes time $O(dr)$ to initialize $M'$.
        \item In every iteration, Line~\ref{line:sub_init_w} takes time $O(d)$ to initialize $w$. Since there are $r$ iterations, the total running time is $O(dr)$.
        \item In every iteration, Line~\ref{line:sub_update_w} takes time $O(1)$ to compute $(\wt{H}_{s+k})^2$ and then takes time $O(d)$ to compute $(w \circ w) / a$. Since there are $r$ iterations, the total running time is $O(dr)$.
        \item In every iteration, Line~\ref{line:sub_update_i} takes time $O(d\log d)$ to sort $w$ (without overwriting) and takes time $O(d)$ to read the indices of top $(1-p)d$ largest entries. Since there are $r$ iterations, the total running time is $O(r(d \log d + d)) = O(r d \log d)$.
        \item In every iteration over $k$, Line~\ref{line:sub_update_m_ki} takes time $O(d)$ to update $M'$. Since there are $r$ iterations, the total running time is $O(dr)$.
    \end{itemize}
    Hence the Procedure \textsc{MaskSelect} takes time
    \begin{align*}
        O(dr + dr + r + dr + rd\log d + dr) = O(rd\log d).
    \end{align*}

    Now, we analyze the running of Line~\ref{line:update_m} over all iterations in Algorithm~\ref{alg:sparse_gpt}. Let $r = B_s$.
    The total number calls to \textsc{MaskSelect} is $d/B_s$. Hence the running over all iterations is
    \begin{align*}
        O((d/B_s) \cdot (B_s d \log d)) = O(d^2 \log d).
    \end{align*}
    Thus we complete the proof.
\end{proof}

Next, we analyze the running time of several key steps of Procedure \textsc{SparseGPT} in Algorithm~\ref{alg:sparse_gpt}. We show that the inverse Hessian can be computed in time $O(d^\omega)$.

\begin{lemma}\label{lem:run_time_h}
    The running time of Line~\ref{line:init_h} in Algorithm~\ref{alg:sparse_gpt} is $O(d^\omega)$.
\end{lemma}
\begin{proof}
    Line~\ref{line:init_h} takes time $O(d^\omega)$ to compute $XX^\top$, takes time $O(d)$ to add $XX^\top$ and $\lambda I_{d\times d}$. Then computing the inverse of a $d \times d$ matrix takes time $O(d^\omega)$. Hence Line~\ref{line:init_h} takes time
    \begin{align*}
        O(d^\omega + d + d^\omega) = O(d^\omega),
    \end{align*}
    where it follows from $\omega \geq 2$.
\end{proof}

Now, we compute the running time of updates of $W$ in the inner iterations.

\begin{lemma}\label{lem:run_time_2}
The running time of Line~\ref{line:update_w_inner} in Algorithm~\ref{alg:sparse_gpt} over all iterations is $O(d^{2+a})$.
\end{lemma}
\begin{proof}
In every iteration, Line~\ref{line:update_w_inner} involves one matrix addition and one matrix multiplication. The size of $E_{*, j-i}$ is $d \times 1$ and the size of $\wt{H}_{j, [j, i+B]}$ is $1 \times B$. Hence multiplying $E_{*, j-i}$ with $\wt{H}_{j, [j, i+B]}$ takes time $O(dB)$. Subtracting $E_{*, j-i}\wt{H}_{j, [j, i+B]}$ from $W_{*, [j, i+B]}$ takes time $O(dB)$. Since there are $d$ iterations, the running time of Line~\ref{line:update_w_inner} in Algorithm~\ref{alg:sparse_gpt} over all iterations is 
\begin{align*}
    O(d \cdot (dB + dB)) = O(d^2B)= O(d^{2+a}),
\end{align*}
where the first step uses basic algebra and the second step is due to $B = d^a$. 
\end{proof}

Finally, we provide the analysis of the running time for updating $W$ in the outer iterations.

\begin{lemma}\label{lem:run_time_1}
The running time of Line~\ref{line:update_w_outter} in Algorithm~\ref{alg:sparse_gpt} over all iterations is $O(d^{1+\omega(1,1,a) - a})$.

\end{lemma}
\begin{proof}
In every iteration, Line~\ref{line:update_w_outter} involves one matrix addition and one matrix multiplication. The size of $E$ is $d \times B$. The size of $\wt{H}_{[i,i+B], [i+B,d]}$ is $B \times (d-(i+B))$. Multiplying $E$ with $\wt{H}_{[i,i+B], [i+B, d]}$ takes time $\Tmat(d, B, d-(i+B)) = \Tmat(d, B, d)$. Subtracting $E\wt{H}_{[i, i+B], [i+B,d]}$ from $W_{*, [i+B,d]}$ takes time $O(dB)$.
Hence, the running time of Line~\ref{line:update_w_outter} over all iterations is
\begin{align*}
\sum_{i=1}^{d/B} O(dB +  \Tmat(d, B, d))
= & ~ \sum_{i=1}^{d/B} O(\Tmat(d, B, d)) \\
= & ~ (d/B) \cdot O(\Tmat(d, B, d)) \\
= & ~ d^{1-a} \cdot O(\Tmat(d,d^a, d)) \\
= & ~ d^{1-a} \cdot O({d^{\omega(1,1,a)}}) \\
= & ~ O(d^{1 + \omega(1,1,a) - a}),
\end{align*}
where the first step is because $\Tmat(d, B, d) \geq \Omega(dB)$, the second step follows from basic algebra, the third step uses $B=d^a$, the fourth step is due to Definition~\ref{def:omega_alpha}, and the last step uses basic algebra.
\end{proof}

\section{Conclusion}\label{sec:conclusion}
We improved the complexity analysis of SparseGPT from $O(d^3)$ to $O(d^{2.53})$, using techniques from fast matrix multiplication and lazy update ideas used in iterative maintenance problems. This tighter bound demonstrates that large language models can be pruned more efficiently than previously thought. Future work could explore further improvements or extensions of these methodologies to other model compression algorithms.

\ifdefined\isarxiv
\section*{Acknowledgement}
Research is partially supported by the National Science Foundation (NSF) Grants 2023239-DMS, CCF-2046710, and Air Force Grant FA9550-18-1-0166.
\else

\fi

\ifdefined\isarxiv
\bibliographystyle{alpha}
\bibliography{ref}
\else
\bibliography{ref}
\bibliographystyle{alpha}
\fi





\end{document}